Two points charges are placed on the x-axis: a – 6 µC charge at x = 0 and a + 5µC charge at x = 200 cm. At
what point(s) in the vicinity of these two charges can a + 5- µC charge be placed to that it experiences no
resultant force?
1
Expert's answer
2020-06-30T18:24:16-0400

Let x be the coordinate of the third charge. The force from the first charge is

"\vec{F_1} = k\dfrac{q_1q_1}{r_{12}^3}\vec{r}_{12}" , "|F_1| = k \cdot\dfrac{5\,\mathrm{\mu C}\cdot 6\,\mathrm{\mu C}}{(x-0)^2} = k\cdot\dfrac{30}{x^2}" and this force is directed from x to 0.

The force from the second charge is

"|F_2| = k \cdot\dfrac{5\,\mathrm{\mu C}\cdot 5\,\mathrm{\mu C}}{(x-2)^2} = k\cdot\dfrac{25}{(x-2)^2}" and this force is directed from 2m to x.


Let the charge be placed between two charges. The net force is 0, so

"\dfrac{-30k}{x^3}x - \dfrac{25k}{(x-2)^3}(x-2) = 0. \;\; \dfrac{-30k}{x^2} - \dfrac{25k}{(x-2)^2} = 0." "\;\;\; \dfrac{6}{x^2} =- \dfrac{5}{(x-2)^2}."

"11x^2-24x+24 = 0." This equation has no roots.


Let the charge be placed left to the first charge, so

"\dfrac{30k}{x^2} - \dfrac{25k}{(x-2)^2} = 0. \;\;\; \dfrac{6}{x^2} =\dfrac{5}{(x-2)^2}." "x^2-24x+24 = 0. \;\; x = 12 \pm2\sqrt{30}" , but none of these roots is less than 0.


Let the charge be placed right to the second charge, so

"-\dfrac{30k}{x^2} -+\dfrac{25k}{(x-2)^2} = 0. \;\;\; \dfrac{6}{x^2} =\dfrac{5}{(x-2)^2}." "x^2-24x+24 = 0. \;\; x = 12 \pm2\sqrt{30}" . We choose the root "x = 12+2\sqrt{30} \approx 22.95\,\mathrm{m}."